OpenStudy (lifeisadangerousgame):

I need help solving y = -2x + 9 and y = 4x - 3 I need to find out what x and y are

OpenStudy (hoblos):

y=y -2x + 9 = 4x - 3 6x=12 x=2 y = 4x - 3=5

OpenStudy (hoblos):

(2,5)

OpenStudy (lifeisadangerousgame):

Thank you :) Do you think you could show me how to solve them separately too?

OpenStudy (hoblos):

separately!! what do you mean?

OpenStudy (lifeisadangerousgame):

like they're different equations one is y = -2x + 9 and the other is y = 4x - 3

OpenStudy (anonymous):

Input Interpretation: y= -2x+9 y= 4x-3 Results: x = 2 and y=5

OpenStudy (lifeisadangerousgame):

but how do I do that Hershey?

OpenStudy (hoblos):

but these are the same equations!!

OpenStudy (anonymous):

^agreed

OpenStudy (anonymous):

@Life what are you talking about ...

OpenStudy (lifeisadangerousgame):

Nevermind, I just don't understand, thanks though :)

OpenStudy (anonymous):

ok....

OpenStudy (hoblos):

it is important that you understand so that you know how to solve such problems which part you are not understanding ?

OpenStudy (anonymous):

yeah..which part do you need explaining for?

OpenStudy (lifeisadangerousgame):

I'm not sure how they're the same question

OpenStudy (hoblos):

same equations

OpenStudy (lifeisadangerousgame):

equations I mean

OpenStudy (hoblos):

hmmmm do you mean why did we assume that y=y ??

OpenStudy (lifeisadangerousgame):

yea

OpenStudy (hoblos):

because that's the way to solve such problems!!

OpenStudy (lifeisadangerousgame):

ohh, I see now, Imma review the lesson again. Thanks :)

OpenStudy (hoblos):

my pleasure XD